jessica makes 15% profit on lunches and 20% profit on dinners. Jessica took in $3000 on tuesday and made a profit of $539.50, how much was spent at lunch?
Answer by Theo(13342)

x = the amount she made on lunches.
y = the amount she made on dinners.

her profit on lunches was .15x
her profit on dinners was .20x

she took in a total of 3000 dollars.
x + y = 3000

her total profit was 539.50
.15x + .20y = 539.50

you have 2 equations that need to be solved simultaneously because the same solution applies to both.

they are:

x + y = 3000
.15x + .20y = 539.50

we'll solve by elimination.

multiply both sides of the second equation by 5 and leave the first equation as is to get:

x + y = 3000
.75x + y = 2697.50

subtract the second equation from the first to get:

.25x= 302.50

divide both sides of this equation by .25 to get x = 1210.

since x + y = 3000, then x = 1210 and y = 3000 - 1210 = 1790.

she made 1210 on lunch and 1790 on dinner for a total of 3000 dollars.

she made .15 * 1210 = 181.50 profit on lunch and .20 * 1790 = 358.00 profit on dinner for a total profit of 181.50 + 358.00 = 539.50.

solution looks good.

the solution is that 1210 dollars was spent on lunch.